OpenStudy (anonymous):

calculus question: find a function f such that the graph of f has a horizontal tangent at (2,0) and f''(x)=2x.

OpenStudy (anonymous):

will you give medal

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can you help me please?

OpenStudy (anonymous):

yeah but do you have a simple pic i can see of it it don't make scene but i can still help

OpenStudy (anonymous):

@Studentc14

OpenStudy (anonymous):

There's no picture it's just words on my homework

OpenStudy (tylerd):

hmm f''=2x so wouldmt f'(x)= x^2

OpenStudy (tylerd):

ah i think i got it

OpenStudy (tylerd):

\[f"(x) =2x\] \[f'(x)=x^2+consant\] \[f'(2)=2^2+constant=0\] constant would then = -4

OpenStudy (tylerd):

hmm

OpenStudy (tylerd):

i think the original function is \[f(x)=\frac{ x^3 }{ 3 }-4\]

OpenStudy (tylerd):

-4x i mean

OpenStudy (tylerd):

ya thats it

OpenStudy (tylerd):

see what i did there?

OpenStudy (tylerd):

\[d(\frac{ x^3 }{ 3 } - 4x = x^2-4\] then the derivative of that is 2x

OpenStudy (anonymous):

ah yes i see! thank you @TylerD

OpenStudy (tylerd):

this is a derivative + integrals problem. and im not to good with integrals but was able to solve it by eyeballing it basically without taking the integral

OpenStudy (anonymous):

yeah i see, thank you! would it also have +c at the end after -4x or no? @TylerD

OpenStudy (tylerd):

no constant was just the placeholder so i know theres something there then solved for it

OpenStudy (tylerd):

actually it says (2,0) so wait a min here

OpenStudy (anonymous):

i mean in the final solution, because couldn't x^3/3 -4x still be shifted up or down and have the same derivatives?

OpenStudy (tylerd):

meh i tink thats right though

OpenStudy (anonymous):

ok thank you

OpenStudy (tylerd):

ya it could be shifted

OpenStudy (tylerd):

it would be shifted up by like +5.333

OpenStudy (anonymous):

ok thanks!

OpenStudy (freckles):

\[f''(x)=2x \\ f'(x)=x^2+b \\ f(x)=\frac{x^3}{3}+bx+c\] so we know f looks like that above we want to find the tangent to f such that there is a horizontal tangent at (2,0) This also means (2,0) belongs to the function f so f'(x)=0 when x=2 so b=4 is right \[f(x)=\frac{x^3}{3}+4x+c \\ \text{ then use (2,0) \to find c } f(2)=\frac{8}{3}+8+c \\0=\frac{8}{3}+8+c \text{ since } f(2)=0 \\ 0=\frac{8}{3}+\frac{24}{3}+c\] then solve for c

OpenStudy (freckles):

oops b=-4

OpenStudy (freckles):

\[f(x)=\frac{x^3}{3}-4x+c \\ f(2)=\frac{8}{3}-8+c \\ 0=\frac{8}{3}-\frac{24}{3}+c \\\]

OpenStudy (anonymous):

thank you!! @freckles